Fermi National Laboratory
 

Operation Summary for February 3, 2010 - February 5, 2010

 

Wednesday Feb. 3

The day shift began with the Tevatron (TeV) in access, with Operations monitoring the stash (334.9E10), with the Antiproton Source stacking (17.8E10), with Linac, Booster, and the Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments accessing their collision halls.

During the day shift, Operations halted stacking at 9:45 AM so Pbar personnel could realign the timing of an Accumulator RF station (ARF1). Operations began the TeV checkout and turn-on at 10:15 AM. Pbar resumed stacking at 11:30 AM. A FIRUS high water alarm for MI-65 posted at 2 PM; operators disabled the NuMI beam and prepared the area for an access. Operations checked the orbits and then put the TeV into HEP shot setup at 2:45 PM. Operations halted stacking again to address ARF1 alignment issues.

During the evening shift, Operations established store 7576 at 4:15 PM with an Average Initial Luminosity (AIL) of 265E30. Pbar resumed stacking at 5 PM. NuMI personnel reported that they had restored power to the tripped sump pumps in MI-65; Operations reestablished the NuMI beam at 7 PM.

Thursday Feb. 4

The midnight shift began with Operations monitoring store 7576 (102.51E30) and the stash (303.1E10), with Pbar stacking (24.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7576 at 4:45 AM. Operations put the TeV into shot setup and established store 7577 at 6:15 AM with an AIL of 273E30.

During the day shift, at 11:30 AM, store 7577 aborted due to an instrumentation fault caused by a focusing sextupole power supply (T:SF). The modulator for BRF16 developed an internal fault. RF technicians disconnected the anode for repairs; it was reconnected later in the shift. EE Support personnel reported at 1:15 PM that they found and replaced a bad cable for T:SF. TeV personnel conducted beam studies for the remainder of the shift.

During the evening shift, Operations put the TeV into shot setup at 5 PM and established store 7580 at 6:10 PM with an AIL of 265E30. Operators continued to work at improving beam positions in Booster and MI. There were several ion pump (IP606D) trips, causing vacuum burps in the MI RF region.

Friday Feb. 5

The midnight shift began with Operations monitoring store 7580 (118.23E30) and the stash (172.1E10), with Pbar stacking (22.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7580 at 6:15 AM and put the TeV into shot setup; both protons and antiprotons had been loaded prior to the end of shift.

A M01 LCW system developed a pump issue delaying the establishment of beam to MTest until Mechanical Support personnel could arrive to investigate. Operations prepared M01 for an access.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/05/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory